Arrange the following elements by core charge. Highest core charge Lowest core charge Answer Bank P S Na Al Cl
Answer #1

Answer :- Core Charge af an elment = number of Protons in the muchos - number of Core electrons finner Shell electros. Ü di n(1) Na: Poral Protors= 11, 2.8.1 no. al Core electrons = 10 - A core change al Na=11-10=1 () P: Total Protors= 15 , 2,8,5 No